write the equation of a line perpendicular to y=-4x+6 and passes through (-12,5)
Answer by Nate(3500)

since the equation is perpendicular to the line of y=-4x+6 you would take the opposite of the recipricol of that slope (1/4)
we need to use the point-slope equation which is: y-(y value)=m(x-(x value))
y-(y value)=m(x-(x value)) --original equation
y-5=(1/4)(x+12) --plug in the points from the ordered pair
y-5=(1/4)x+3 --simplify
y=(1/4)x+8 --add 5 to each side